OpenStudy (anonymous):

The cost of manufacturing x cases of fillfluff is C dollars (in thousands), where C(x) = 4 x + 4 sqrt(x) + 5 Monthly production at t weeks is expected to be x = 6400 + 90 t Write an equation to express rate of change of cost with respect to time: (dC)/(dt) =

OpenStudy (anonymous):

i found equation to express manufacturing cost as a function of time: C(t) =25600+360t+4sqrt(6400+90t)+5 and the equation to express marginal cost: (dC)/(dx) =4+2/sqrt(x)

OpenStudy (anonymous):

i agree with c(t) but i think the derivative is 4 + 8x did you take the derivative of initial c(x) ?

OpenStudy (anonymous):

yes i took the deriv and plugged x into the values

OpenStudy (anonymous):

i see what you didd - go back to c(t ) that sqrt is not correct, sorry c(t) = 25600 + 36t + 6400^2 + 43200t + 81t^2 +5 c't = 36 + 43200 + 162t c't = 162t + 43236 your thoughts?